# Cascade-driven opinion dynamics on social networks

### v2

@@ V-B3 Polarization
Figure 4 illustrates the polarization outcomes across different settings. As expected, the Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3). Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization. Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude. This behavior can be explained by Table II, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity. Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging. In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
@@ V-C1 Dataset description and experimental setting
In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [50], a Facebook graph [51], and a synthetic Barabási-Albert network [52]. The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot. The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes. Figure 5 shows visualizations of all three graphs.
Figure 5: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts, which are kept fixed across all simulations. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios. In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table II). To explore other potential values, we referenced typical ranges established in the literature. Hodas et al. in [53] estimated that the retweet probability for each tweet seen is around 0.15% to 0.4%. Based on this, we modelled a moderate diffusion scenario with $\theta=0.1$, and for maximum diffusion, we used $\theta=0.5$. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 independent simulation runs.
@@ V-C2 Opinion evolution under FJC vs FJ
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. Due to space constraints, we will present just $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, the best local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}-P_{3}$ polarization, respectively. Please refer to the Supplemental Information (SI) for the full results. In Figures 6-8, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
@@
Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 6 and 7. In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared. Notably, in the Karate network, results remain almost unchanged across different $\theta$ values, as its small size allows most nodes to be reached by the cascade even when propagation is limited. Interestingly, central nodes in the FJC model maintain stable behavior even with tighter propagation. This resilience is an intrinsic property of the model, stemming from the structured, sequential nature of its cascades. In fact, highly connected nodes are typically exposed to content early, when active sources are few, making them less susceptible to external influence.
@@ VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We validated the FJC model, utilising a real cascade dataset. Our results demonstrate that the FJC model can reliably approximate final opinions accurately, even when many of the underlying assumptions of the model are relaxed or only partially met. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution.

### v3

@@ V-B3 Polarization
Figure 4 illustrates the polarization outcomes across different settings. The Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3). Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization. Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude. This behavior can be explained by Table III, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity. Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging. In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics for Nepal top cascades. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
@@ V-C1 Dataset description and experimental setting
In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [55], a Facebook graph [34], and a synthetic Barabási-Albert network [4]. The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot. The Barabasi-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes. Please refer to the SI for a graphic visualization of the graphs.
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts and kept them fixed across all simulations, in order to control for variability due to source selection. The properties of the selected seed nodes and their centrality profiles are analyzed in the SI, where we show that this choice does not introduce systematic bias. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios. In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table III). To explore other potential values, we referenced typical ranges established in the literature. Hodas et al. in [32] estimated that the retweet probability for each tweet seen is around 0.15% to 0.4%. Based on this, we modelled a moderate diffusion scenario with $\theta=0.1$, and for maximum diffusion, we used $\theta=0.5$. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 independent simulation runs.
@@ V-C2 Opinion evolution under FJC vs FJ
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. Due to space constraints, we will present $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, which are the local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}\text{-}P_{3}$ polarization, respectively, in the largest subdomain. Please refer to the SI for the full results. In Figures 5-7, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
@@
Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 5 and 6. In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared. Notably, in the Karate network, results remain almost unchanged across different $\theta$ values, as its small size allows most nodes to be reached by the cascade even when propagation is limited. Interestingly, central nodes in the FJC model maintain stable behavior even with tighter propagation. This resilience is an intrinsic property of the model, stemming from the structured, sequential nature of its cascades. In fact, highly connected nodes are typically exposed to content early, when active sources are few, making them less susceptible to external influence.
@@ VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin-Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We validated the FJC model, utilising a real cascade dataset. Our results demonstrate that the FJC model can reliably approximate final opinions accurately, even when many of the underlying assumptions of the model are relaxed or only partially met. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution.